Solve by elimination
How do I solve 4x-3y=39
-3x-4y=23
I have worked this problem over and over please solve this for me.

Answer by CharlesG2(834)

"Solve by elimination
How do I solve 4x-3y=39
-3x-4y=23
I have worked this problem over and over please solve this for me."
4x - 3y = 39 (multiply by 3)
-3x - 4y = 23 (multiply by 4)
12x - 9y = 117
-12x - 16y = 92 (add 1st to 2nd)
12x - 9y = 117
0x - 25y = 209 (solve this one for y)
y = 209/(-25) = -209/25 = -8 9/25 = -8.36
plug in this y into 1st equation and solve for x
4x - 3(-8.36) = 39
4x + 25.08 = 39
4x = 13.92
x = 3.48
-3x-4y=23 (plug x and y into this one to check)
-3(3.48) - 4(-8.36) =? 23
-10.44 + 33.44 = 23 (works)